Solve each equation. Identify each equation as a conditional equation, an inconsistent equation, or an identity.
1/2a - 5 = 1/3a - 1

Fill in the tables using the given formulas
y = -5x + 10
________________________
x y
_______________________
-20
-10
0
10
______________________

Answer by zeynep(43)

1/2a - 5 = 1/3a - 1
let's get rid of those fractions by multiplying both sides by the LCD of 6;
6.(1/2a - 5) =6.( 1/3a - 1)
3a-30=2a-6 add additive inverse of 2a to both sides;
3a-30-2a=2a-6-2a
a-30=-6 add additive inverse of -30 to both sides;
a-30+30=-6+30
a=24
This equation is a conditional equation. A conditional equation is an equation that is not true for all real numbers but has at leat one real number solution.
I could not understand your second question. Are the given values for x?